• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name M. Liang

    Log entry time 00:02:02 on September 23,1999

    Entry number 26996

    keyword=Swing shift activities

    18:44 Arc energy done. Move C target in position. Requested 10 muA. No raster.
    18:50 Start Run 1747. Set ps1=ps2=ps3=ps5=1, ps4=100. TS3 is about 800 Hz,
    TS1 is about 200Hz. Will take data for 5 minutes.
    Beam off most of the time.
    19:02 Stop run 1747, start a new run 1748. Beam stable most of the time.
    19:08 stop run 1748. Ask MCC to mask target movement. Move to dummy target.
    Request raster on (2*2). Same prescale factor as before.
    Start run 1749. Run for 10 minutes.
    Checked spot size: looks good. 2x2.
    Beam tripped, MCC thought there was target movement.
    19:30 stop run 1749. Switch to 15cm LH2 target.
    19:34 Asked for 10 muA. TS3 rate is about 3kHz. Reduce beam current to 3 muA.
    Keep the same prescale factor. Start run 1750. 10 minutes.
    Start Run 1751. for another 10 minutes.
    20:03 Moved radiator in 6%.
    20:06-20:16 Run 1752. ~1.5 M total
    20:17-20:28 Run 1753. ~1.93 M total
    20:28-20:40 Run 1754. ~1.68 M total
    20:41-21: 09 Run 1755. Changed ps1/ps2/ps3/ps4/ps5 to 4/2/1/400/1. File split.
    21:09-21:22 Run 1756. File split again.
    21:22 changed to dummy target.
    21:30-21:47 Run 1757.
    21:47 changed Earm,Harm configurations to Pe=0.860 GeV/c, theta_e=45.28deg,
    Ph = 0.8398, theta_h=46.709deg.
    Asked beam current to crease to 10 muA.
    22:15-22:27 Run 1758. <=== need to be redone, forget 3" Carbon door
    22:27 Added 3" Carbon door.
    22:33-22:46 Run 1759. with 1.5"+3" Carbon door.
    22:46 Switched to LH2 target. TS3 rate is very high. ~6.6kHz at 10 muA. Asking for
    3 muA now.
    22:53-23:10 Run 1760. ps1/ps2/ps3/ps4/ps5: 4/10/1/1000/1. 1.8M total
    23:11-23:31 Run 1761. 1.9M total
    23:31-23:52 Run 1762
    23:52 Moved radiator out, increase current to 4 muA.
    23:57 Start run 1763.